OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. M(–2, 1) and O(–3, 2)

OpenStudy (lgbasallote):

hint: \[\huge m_x = \frac{x_1 + x_2}2\] \[\huge m_y = \frac{y_1 + y_2}2\] does that help?

OpenStudy (anonymous):

yes it does